OpenStudy (anonymous):

can you help with this one @jim_thompson5910

OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (1, 6), what quadratic function is created? f(x) = one eighth (x - 1)2 - 2 f(x) = -one eighth (x + 1)2 - 2 f(x) = -one sixteenth (x + 1)2 - 2 f(x) = one sixteenth (x - 1)2 + 2

OpenStudy (anonymous):

please send a picture of the problem

OpenStudy (anonymous):

and if u need sustainable aid contact me the,

OpenStudy (anonymous):

then